# GeoDANO: Geometric VLM with Domain Agnostic Vision Encoder

### v1

@@ C.1 Domain adaptation data
Figure 6: Examples of diagram pairs curated for domain adaptation. For each row, the first diagram is from the target domain, and the remaining diagrams are from the source domain. To generate source domain diagrams, we translate the target diagram by our diagram generator with the textual description of the target image.
C.2 Training details
@@
We start from OpenCLIP (Radford et al., 2021), a pre-trained model where the architecture is ViT-L/14 with image resolution $336\times 336$. To train OpenCLIP, we use total of 200,000 diagram-caption pairs generated with our synthetic data engine. For the domain adaptation to GeoQA and Geometry3K datasets, we randomly sample 50 diagrams and translate the diagram and caption styles following the procedure described in Section 4.2. Finally, GeoCLIP is fine-tuned via Eq. 2. We name the GeoQA and Geometry3K adopted GeoCLIP as GeoCLIP-DA.
[…]
Appendix D GeoDANO
D.1 Modification of training data
Our fine-tuning strategy differs slightly from previous works (Chen et al., 2022; Zhang et al., 2023; Xia et al., 2024). Here, we clarify the difference between our approach and previous approaches. In previous works, the VLM is trained to produce the solution program given diagram and problem description as shown in Fig. 1. An interesting observation from GeoQA and PGPS9K datasets is that the numerical measurements, such as angles, lengths, and volumes, are not written in the problem description but given as additional conditions, and the numerals are substituted as a variable in the problem description as shown in Fig. 1(a). Therefore, the VLM only needs to produce the solution program without having optical character recognition (OCR) from the diagram. The variables are automatically substituted by the actual numbers when the program is executed. Therefore, the vision encoders do not need to learn OCR from the image.
However, this approach cannot be generalized to a wider class of problems where the numerals are embedded in the diagram instead of written in the problem description. Some variants of MathVerse, such as the vision-dominant problems, fall into this category as well. To incorporate OCR into the solution of the problem, we modify some problem statements in the training set, such that the numerical measurements are only shown in the diagram and not in the statements. We further modify the solution problem so that the solution contains OCR results as a part of the final output. Finally, we unify the language of the solution programs used in GeoQA and PGPS9K by converting GeoQA programs into PGPS9K format. The unification makes the output of VLM consistent since both datasets use different types of formal languages.
Fig. 7 shows examples of the modified input pairs and solutions, where the first problem statement does not have numerical measurements and the OCR results are in the part of the output solution program.
@@ D.1 Modification of training data
Figure 7: Examples of the training data for GeoDANO. While previous PGPS models require the only to predict the solution steps and assume the numerical values are explicitly given, GeoDANO is trained to predict both the solution steps and the numerical values in the diagram and text.
D.2 Training details
@@
With the modified training data, we apply supervised fine-tuning on the VLM, i.e., the gradient only flows through the prediction of numerical values and solution steps, not the diagram and text.
[…]

### v2

@@ D.1 Details of the translation process
Figure A4: Examples of diagram pairs curated for domain adaptation. For each row, the first diagram is from the target domain, and the remaining diagrams are from the source domain. To generate source domain diagrams, we translate the target diagram by our diagram generator with the textual description of the target image.
D.2 Domain adaptation data
We adopt GeoCLIP to the two PGPS benchmarks: GeoQA (Chen et al., 2021) and PGPS9K (Zhang et al., 2023). For PGPS9K, we use the Geometry3K split. Fig.˜A4 shows the pairs used to adapt the domain of GeoCLIP.
D.3 Training details
@@
We start from OpenCLIP (Radford et al., 2021), a pre-trained model where the architecture is ViT-L/14 with image resolution $336\times 336$. To train OpenCLIP, we use total of 200,000 diagram-caption pairs generated with our synthetic data engine. For the domain adaptation to GeoQA and Geometry3K datasets, we randomly sample 50 diagrams and translate the diagram and caption styles following the procedure described in Section˜4.2. Finally, GeoCLIP is fine-tuned via Eq.˜2. We name the GeoQA and Geometry3K adopted GeoCLIP as GeoCLIP-DA.
We set the batch size for the source domain diagram-caption pairs to 256. For the domain adaptation parts, i.e., applying CLIP on the diagram-caption pairs and the diagram pairs of target domains, we vary the batch size to 32. We set weight decay to 0.2. We optimize for 50 epochs using Adam optimizer (Kingma, 2014) and a cosine annealing scheduler with 2,000 warmup steps, and the maximum learning rate is set to be $1\text{e-}4$. We train the model with eight RTX3090 GPUs for approximately 24 hours.
D.4 Visualization of GeoCLIP-DA embeddings
We compare the embeddings between GeoCLIP-DA and OpenCLIP in Fig.˜A5.
Figure A5: Visualization of OpenCLIP and GeoCLIP-DA embeddings. The orange, green, and blue dots represent PGPS9K, GeoQA, and synthetic diagrams, respectively. In the top row, the three diagrams on the left and right are those with the highest cosine similarities to the center under OpenCLIP and GeoCLIP-DA, respectively.
Appendix E GeoDANO
@@
Figure A6: Examples of the training data for GeoDANO. While previous PGPS models require only predicting the solution steps and assuming the numerical values are explicitly given, GeoDANO is trained to predict both the solution steps and the numerical values in the diagram and text.
E.1 Training details
Architectural details.
@@ E.1 Training details
Training approach.
With the modified training data, we apply supervised fine-tuning on the VLM, i.e., the gradient only flows through the prediction of numerical values and solution steps, not the diagram and text. During the training of GeoDANO, the parameters of the vision encoder, i.e., GeoCLIP-DA, are frozen and remain unchanged. The projection layer, which maps visual embeddings to language model inputs, is randomly initialized and trained from scratch simultaneously with the language model.
Hyper-parameters.
We train the VLM with AdamW optimizer (Loshchilov and Hutter, 2019) and cosine annealing scheduler with warmup ratio 0.03 and maximum learning rate $1\text{e-}5$. We use LoRA (Hu et al., 2022) with rank 128. We set the batch size to 16 and train with 5 epochs. We train the VLM with four A100-80GB GPUs for approximately 24 hours.